OpenStudy (anonymous):

What are the sine, cosine, and tangent of Θ = 3 pi over 4 radians?

OpenStudy (anonymous):



OpenStudy (luigi0210):

\(\frac{3\pi}{4}=135^○\)

OpenStudy (anonymous):

where do you get the 135 from?

OpenStudy (luigi0210):

Just \(\frac{\pi}{4}\) is 45 degrees

OpenStudy (anonymous):

sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = -1 sin Θ = negative square root 2 over 2; cos Θ = square root 2 over 2; tan Θ = 1 sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = 1 sin Θ = negative square root 2 over 2; cos Θ = square root 2 over 2; tan Θ = -1 Are my answer choices btw, if that helps.

OpenStudy (anonymous):

okay, so it's 3 times 45 basically. Right?

OpenStudy (anonymous):

http://puu.sh/aTmll/3df5f39361.png Thats the answer choices again, probably easier to tell like this.

OpenStudy (anonymous):

What do I do with the 3x/4 = 135 exactly?

OpenStudy (luigi0210):

Since 3pi/4 is a multiple of 45, just in a different quad, use the 45-45-90 triangle

OpenStudy (anonymous):

What does the little circle with an H in it mean btw? Hypotenuse?

OpenStudy (luigi0210):

It would look something like this: